# Scalability Challenges in Variational Quantum Optimization under Stochastic Noise

### v1

@@ III.5 Measures of Solvability
should be considered the solution candidate of the VQA. Note that in a real experiment, this candidate would be chosen based on the basis state with the highest measured shot count. However, since we are working with simulated results, we can directly use the exact definition.
@@
where $N$ is the total number of optimization runs. This probability reflects the fraction of runs among all experiments that yielded a solution within the specified threshold $t$. In the results presented in Section IV, we denote the probability of finding the exact optimal solutions (${t=1}$) as $\hat{p}_{\mathrm{opt}}$, while near-optimal solutions within a $5\%$ and $1\%$ margin of the optimal are monitored as $\hat{p}_{95\%}$ and $\hat{p}_{99\%}$, respectively.
@@
IV Results of Scaling Analysis
@@
IV.1 Solvability under Additive Gaussian Noise
@@
IV.2 Solvability Depending on System Size
@@
Figure 7: Noise resilience parameters as a function of system size. (a) The point of steepest descent $\sigma^{*}$, (b) slope at steepest descent $m^{*}$, and (c) point of resilience $\sigma_{\mathrm{res}}$ are shown for increasing system sizes $n$, derived from the $\tanh$-fits of $\hat{p}_{\mathrm{opt}}(\sigma)$ – the exact solvability – of various optimizers. All plots use logarithmic y-axes for better visualization.
Beyond ranking the robustness of the optimizers, we examine the overall scaling behavior of noise resilience metrics, which appear to follow consistent trends across different optimizers. Due to the strong similarity in results between $\hat{p}_{\mathrm{opt}}$ and $\hat{p}_{99\%}$, Fig. 7 focuses on exact solvability only. Lower success thresholds (e.g., $\hat{p}_{95\%}$) exhibit slight deviations, which are discussed in Section VI and Appendix E.
The point of steepest descent $\sigma^{*}$ and the point of resilience $\sigma_{\mathrm{res}}$ should, in principle, exhibit similar scaling behavior, as both represent specific x-positions of the same $\tanh$-curve. All three metrics – including the slope of the descent’s tangent line $m^{*}$ – show a notable decline with increasing system size. Given the limited experimental data – constrained by the computational cost of simulating larger system sizes – any conclusions about the exact scaling behavior or asymptotic trends require cautious analysis. To better understand the nature of these decay trends, we therefore fit three different decay functions to the boundary line $\sigma^{*}(n)$:
@@
The results of these fits are illustrated in Fig. 8. Visually, the power-law fit appears to align best with the majority of optimizer curves. However, the fit residuals do not necessarily support this hypothesis. They are given by
@@
Despite uncertainties regarding the exact form of decay, one clear conclusion emerges: noise resilience declines rapidly across all tested methods. This decline is intuitively attributed to the corresponding shrinkage of loss function variance (cf. Eq. 14). However, even after normalizing the noise level $\sigma$ to account for this effect (cf. Section V), this decline persists. This suggests that loss variance decay cannot fully explain the observed drop in noise resilience among classical optimizers.
@@ V.1 Hardware Development
V.2 The Curse of Finite Sampling
In the case of fault-tolerant hardware, the only persistent source of stochastic noise is produced by finite measurement sampling. Independent of whether VQAs will be used with fault-tolerant systems, this can be considered the ideal setting for these algorithms. This section, therefore, deals with the question of how the noise thresholds for problem solvability identified in Section IV translate to the required number of measurement shots and what implications their scaling behavior has for achieving a practical quantum speed-up. In the following, we consolidate all previous findings, adopting the most optimistic scenarios.
Based on the results of Section IV.2, the decline in an optimizer’s noise resilience as the system size $n$ increases, denoted by $\sigma^{*}(n)$, may follow an exponential, logarithmic, or power-law trend (cf. Fig. 8). To convert these resilience thresholds of the most noise-resilient optimizer, NGD, into the general measure of the RAE (34), they need to be normalized by the variance of the corresponding loss function (cf. Eq. 35), as obtained for BENQO. The most optimistic scaling for the loss variance, which also provided a good fit, followed a power-law decay  [^4] Using an exponential fit for the loss variance – consistent with a barren plateau assumption – would imply an increase in the noise resilience for system sizes $n\geq 61$ in the power-law case (31), and $n\geq 27$ for logarithmic decay (32). While this would not change the overall implications of Section V.2, this scenario is unlikely.. Using the fit parameters from Fig. 4, the maximum RAE $\epsilon^{*}(n)$ that an optimizer can tolerate for a given system size $n$ computes to
@@ V.2 The Curse of Finite Sampling
To map this to measurement shot requirements, we first analyze the error introduced by finite sampling in evaluating a loss function, denoted as $\epsilon_{\mathrm{FS}}$. Fig. 10 illustrates the behavior of the RAE (34), computed for three VQAs (cf. Section III.2), including BENQO, as a function of the number of shots $n_{\mathrm{shots}}$ (at a fixed system size) and the system size $n$ (at a fixed number of shots). As expected from a multinomial probability distribution, the error decreases proportionally to $1/\sqrt{n_{\mathrm{shots}}}$ when $n$ is fixed. Conversely, when $n_{\mathrm{shots}}$ is kept constant, the error grows exponentially, reflecting the increasing range of loss values alongside a simultaneously shrinking variance as the system size increases. For an analytical derivation of this scaling behavior, see Appendix C (cf. [52]).
@@
The figure clearly illustrates that if the boundary line identified in Section IV.2 was to follow an exponential decay, the number of required measurement shots would already grow faster than the number of brute-force trials needed to solve the problem classically. This strongly suggests that quantum advantage for optimization tasks under these conditions is fundamentally unattainable. In the alternative scenarios of power-law and logarithmic scaling, a small window of potential quantum advantage remains for problem sizes $n\geq 20$ (logarithmic boundary) and $n\geq 25$ (power-law boundary), where the shot requirement still falls below the critical threshold of quantum disadvantage. However, these potential advantages are only theoretical, as they are based purely on resource scaling and do not yet translate into practical runtime improvements.
@@
VI Validation and Discussion
A common concern in empirical studies like ours is whether the observed results are influenced by specific choices in the experimental setup (cf. Fig. 3). Factors such as problem formulation, quantum ansatz, noise model, optimizer selection, and performance metrics could, in principle, affect the conclusions drawn. One may, therefore, argue that the critical behavior identified in this work merely arises from unlucky experimental choices rather than inherent scalability limitations of variational quantum optimization. While each of our decisions was carefully justified to favor best-case scenarios and ensure a fair and optimistic evaluation, this section provides additional validation and discussion to support the reliability and generality of our findings.
VI.1 Independent Confirmation of Scaling Trends
A study by Scriva et al. [17] examined shot noise effects using a VQE ansatz on an Ising chain problem and found that the success probability for a fixed system size follows the same sigmoidal pattern as observed in our experiments. Their results also demonstrated that the optimal number of function calls required to reach a fixed success probability scales exponentially with system size, reinforcing our findings in Section V.2.
@@ VI.1 Independent Confirmation of Scaling Trends
Despite the differences in quantum algorithms, noise models, classical optimizer selection, and problem types, both studies independently observed the same key trends as those presented in this work. This strongly supports the generality of our findings, reinforcing that the scalability limitations of VQAs under noise are most likely no artifacts of specific experimental choices but rather represent fundamental constraints of the variational quantum optimization paradigm.
VI.2 Validation with a VQE Ansatz
We repeat our experiments to further validate our conclusions using a different quantum loss function: the two-local shallow VQE ansatz described in Section III.2. The results for the NGD optimizer are summarized in Fig. 12, closely matching those obtained using BENQOs. This agreement is not only qualitative but also quantitative. Normalizing the noise resilience threshold $\sigma^{*}(n)$ by the variance of the respective loss functions (cf. Fig. 4) – thus mapping both results onto the RAE metric (34) – would likely reveal an even stronger correspondence between the two ansatzes.
Figure 12: Validation of NGD solvability scaling with VQE. (a) Contour plot of the empirical NGD solvability experiments (cf. Fig. 5), this time using VQE as the underlying quantum loss function. (b) Scaling of the success probability $\hat{p}_{\mathrm{opt}}$ across different noise levels for fixed system sizes, fitted with sigmoidal theory curves as in Fig. 6. Blue vertical lines indicate the points of steepest descent $\sigma^{*}$ (27), obtained from $tanh$-fits (cf. Eq. 25). (c) The main noise resilience parameter $\sigma^{*}$ is plotted against increasing system sizes and directly compared to the noise threshold obtained for BENQO (cf. Section IV.2). The curves exhibit a close agreement.
To directly validate the corresponding resource scaling reported in Section V.2, we conduct additional experiments with the VQE algorithm, this time introducing noise solely through finite sampling rather than the additive Gaussian noise model. The shallowness of the VQE ansatz enables simulations with larger system sizes than in previous experiments, extending the range to $n\in[3,13]$. The following results focus on the COBYLA solver, as it demonstrated the highest resource efficiency in terms of quantum function calls and mean runtime (cf. Appendix D).
@@ VI.2 Validation with a VQE Ansatz
VI.3 Reconsidering the Success Metric
One final point of discussion of our study is the choice of success metric. In this study, we measured the sample success probability of finding a solution within a specific relative threshold $t$ of the optimal energy (cf. Eq. 23), with thresholds set at $t=100\%$ (exact optimality), $99\%$, and $95\%$. However, most of the scalability issues discussed are derived from the probability of finding the optimal solution $\hat{p}_{\mathrm{opt}}$. One could argue that the primary goal of quantum optimization, particularly for NP-hard problems, is not necessarily to find the absolute global optimum but rather to generate high-quality approximations that outperform classical solutions.
Our experimental results indicate that the scaling behavior of $\hat{p}_{99\%}$ and $\hat{p}_{95\%}$ differs slightly from that of the exact solution probability. In particular, the further the relative energy threshold $t$ deviates from $100\%$, the more its scaling diverges from the solvability scaling discussed earlier. In some cases – especially in the absence of noise – the near-optimal success probability $\hat{p}_{95\%}$ even increases with system size (cf. Fig. 5), suggesting that as the system grows, optimizers tend to converge to local minima that lie within a close vicinity of the global optimum.
This trend can be explained by examining the distribution of loss values across all possible solution states (cf. Appendix E). In our QUBO formulation, feasible solutions are enforced via quadratic penalty terms (cf. Appendix B), which can – depending on the chosen penalty factor – significantly expand the loss value range in one direction, often leading to $|C_{\max}|\gg|C_{\min}|$. Due to the chosen normalization, the maximum loss value remains 1, while the majority of loss values concentrate (possibly exponentially) near zero (cf. Section III.2), leading to an asymmetry in the loss value distribution. As the system size increases, this effect intensifies, effectively pushing most local minima closer to the global optimum. Consequently, the approximation ratio in Eq. 22 approaches 1, shifting the threshold for identifying feasible solutions to much higher energy levels – sometimes even beyond the $99\%$ threshold used in this study [91, 76]. This explains why approximate solutions remain relatively accessible, even when finding the exact optimum becomes increasingly difficult.
@@ VI.3 Reconsidering the Success Metric
VII Conclusions and Outlook
@@
The potential of variational quantum algorithms for optimization fundamentally depends on their ability to handle stochastic noise, which arises both from current hardware imperfections and finite sampling – the latter persisting even in fault-tolerant quantum systems. While prior analytical studies have examined the impact of noise on VQAs [18, 15, 80, 16, 17], few have systematically demonstrated how this noise effectively affects their scalability to practically relevant problem sizes. This work provides a rigorous numerical investigation into how classical optimizers – the main computational engine of VQAs – respond to additive Gaussian noise as the size of quantum loss functions scales up. Our findings, validated through extensive testing, empirically demonstrated that noise fundamentally limits the solvability of variational quantum ansatzes, even in a fault-tolerant setting.
@@
We find, that for fixed system sizes, the probability of finding high-quality solutions – including those within $1\%$ and $5\%$ of the optimal energy – decreases in a sigmoidal pattern with respect to the logarithmic noise level. This reveals a sharp performance drop at a critical noise threshold $\sigma^{*}$, which varies significantly across the different optimizers. Using $\tanh$-fits, we extracted these thresholds for system sizes $n\in[3,10]$, suggesting a common scaling trend across all optimizers: their noise resilience declines rapidly, potentially exponentially, as system size increases. Notably, this decline is steeper than what can be explained by barren plateaus (loss concentration), confirming that larger systems inherently present a more challenging loss landscape.
@@
While it is well known that current hardware error rates far exceed what is required for classically relevant problem sizes, our results highlight an even stricter limitation imposed by finite sampling errors. Extrapolating the observed decline in noise resilience showed that, to achieve the loss accuracy necessary for problem solvability, the required number of measurement shots scales prohibitively bad with system size. This leaves minimal room before exceeding the computational resources of classical brute-force sampling, which is already impractical. These findings cast serious doubt on the feasibility of achieving meaningful quantum advantage with VQAs for large-scale optimization problems.
@@
We therefore advocate for a shift in research focus – away from introducing yet new heuristic hybrid algorithms with idealized testing, and toward strategies that directly address the fundamental challenges demonstrated in this work. If quantum utility [93] in optimization remains the goal, especially in the near term, future research should prioritize alternative algorithmic frameworks (and possibly different problem domains) that can provably circumvent these barriers.
@@
One promising avenue for this is improved problem encoding, which aims to reduce the number of required qubits and thereby mitigate some of the scaling issues discussed here [94, 95, 96, 97, 98]. However, while these methods offer potential efficiency gains, they often lack both theoretical performance guarantees and large-scale empirical validation in noise-exposed settings. Moreover, given that the scalability challenges demonstrated here may be inherent in the hybrid nature of VQAs, which rely on classical optimization, it may be more promising to explore fully quantum alternatives outside the variational paradigm. Recent works [99, 31] have begun such investigations, and a similar shift away from hybrid frameworks can be seen in the quantum machine learning community [100]. Additionally, this study has not yet considered alternative computational models, such as Quantum Boson Sampling [101, 102, 103] or Quantum Annealing [104, 105], which may warrant further investigation.
@@
Ultimately, our work provides a general and robust framework for rigorously investigating the scalability and limitations of not only VQAs but also other quantum algorithms for optimization. We encourage researchers to apply and extend our methodology to more promising alternative setups – whether by testing different quantum ansatzes, optimizers, problem instances, or hardware platforms. If future studies uncover scenarios where problem solvability scales more favorably with system size and noise levels, this would be an exciting breakthrough. However, if the scaling behavior observed here proves to be universal and inherent to the classical optimization problem inherent in hybrid architectures like VQAs, it would strongly suggest that quantum advantage in variational quantum optimization is not realistic. In fact, our results indicate that the challenges facing VQAs – both in the NISQ and fault-tolerant era – may be far more fundamental than what can be resolved through hardware improvements, better classical optimizers, or ansatz tuning alone.

### v2

@@ IV.5 Measures of Solvability
should be considered the solution candidate of the VQA. This choice is motivated by the practical perspective that real-world users aim to directly sample a high-quality solution, rather than undertake the impractical effort of searching for one within a delocalized quantum output distribution. Note that in a real experiment, $|\psi_{\text{max}}\rangle$ would be chosen based on the basis state with the highest measured shot count. However, since we are working with simulated results, we can use the exact definition.
@@
where $N$ is the total number of optimization runs. This probability reflects the fraction of runs among all experiments that yielded a solution within the specified threshold $t$. In the results presented in Section V, we denote the probability of finding the exact optimal solutions (${t=1}$) as $\hat{p}_{\mathrm{opt}}$, while near-optimal solutions within a $5\%$ and $1\%$ margin of the optimal are monitored as $\hat{p}_{95\%}$ and $\hat{p}_{99\%}$, respectively.
@@
V Results of Scaling Analysis
@@
V.1 Solvability under Additive Gaussian Noise
@@
V.2 Solvability Depending on System Size
@@
Figure 7: Noise resilience parameters as a function of system size. (a) The point of steepest descent $\sigma^{*}$, (b) slope at steepest descent $m^{*}$, and (c) point of resilience $\sigma_{\mathrm{res}}$ are shown for increasing system sizes $n$, derived from the $\tanh$-fits of $\hat{p}_{\mathrm{opt}}(\sigma)$ (solid line) and $\hat{p}_{95\%}(\sigma)$ (dashed line), for exact and approximate solvability of various optimizers, respectively. All plots use logarithmic y-axes for better visualization.
Beyond ranking the robustness of the optimizers, we examine the overall scaling behavior of noise resilience metrics, which appear to follow consistent trends across different optimizers. Due to the strong similarity in results between $\hat{p}_{\mathrm{opt}}$ and $\hat{p}_{99\%}$, Fig. 7 focuses on exact solvability ($\hat{p}_{\mathrm{opt}}$) and $\hat{p}_{95\%}$ for approximate solvability.
The point of steepest descent $\sigma^{*}$ and the point of resilience $\sigma_{\mathrm{res}}$ should, in principle, exhibit similar scaling behavior, as both correspond to specific positions along the same $\tanh$-curve. All three metrics, including the slope of the descent’s tangent line $m^{*}$, show a notable decline with increasing system size. While this overall trend is consistent across both exact and approximate solvability, the lower success threshold $\hat{p}_{95\%}$ intuitively exhibits a slightly softer decay compared to $\hat{p}_{\mathrm{opt}}$. However, since the practical usefulness of the approximate solvability metric is subject to limitations (see discussion in Section VII.3 and Appendix E), the subsequent analysis focuses on the scaling behavior of exact solvability. We emphasize, however, that the analysis was conducted for both thresholds, and full data is reported for each (see Appendix F). The overall qualitative conclusions remain unchanged, with only minor quantitative differences in the experimental curves.
Given the limited experimental data – constrained by the computational cost of simulating larger system sizes – any conclusions about the exact scaling behaviors or asymptotic trends require cautious analysis. To better understand the nature of the observable decay trends in Fig. 7, we therefore fit three different decay functions to the boundary line $\sigma^{*}(n)$:
@@
The results of these fits are illustrated in Fig. 8 for $\hat{p}_{\mathrm{opt}}$. Visually, the power-law fit appears to align best with the majority of optimizer curves. However, the fit residuals do not necessarily support this hypothesis. They are given by
@@
Despite uncertainties regarding the exact form of decay, one clear conclusion emerges: noise resilience declines rapidly across all tested methods. This decline is intuitively attributed to the corresponding shrinkage of loss function variance (cf. Eq. 14). However, even after normalizing the noise level $\sigma$ to account for this effect (cf. Section VI), this decline persists. This suggests that loss variance decay cannot fully explain the observed drop in noise resilience among classical optimizers.
@@ VI.1 Hardware Development
VI.2 The Curse of Finite Sampling
In the case of fault-tolerant hardware, the only persistent source of stochastic noise is produced by finite measurement sampling. Independent of whether VQAs will be used with fault-tolerant systems, this can be considered the ideal setting for these algorithms. This section, therefore, deals with the question of how the noise thresholds for problem solvability identified in Section V translate to the required number of measurement shots and what implications their scaling behavior has for achieving a practical quantum speed-up. In the following, we consolidate all previous findings, adopting the most optimistic scenarios.
Based on the results of Section V.2, the decline in an optimizer’s noise resilience as the system size $n$ increases, denoted by $\sigma^{*}(n)$, may follow an exponential, logarithmic, or power-law trend (cf. Fig. 8). To convert these resilience thresholds of the most noise-resilient optimizer, NGD, into the general measure of the RAE (34), they need to be normalized by the variance of the corresponding loss function (cf. Eq. 35), as obtained for BENQO. The most optimistic scaling for the loss variance, which also provided a good fit, followed a power-law decay  [^4] Using an exponential fit for the loss variance – consistent with a barren plateau assumption – would imply an increase in the noise resilience for system sizes $n\geq 61$ in the power-law case (31), and $n\geq 27$ for logarithmic decay (32). While this would not change the overall implications of Section VI.2, this scenario is unlikely.. Using the fit parameters from Fig. 4, the maximum RAE $\epsilon^{*}(n)$ that an optimizer can tolerate for a given system size $n$ computes to
@@ VI.2 The Curse of Finite Sampling
To map this to measurement shot requirements, we first analyze the error introduced by finite sampling in evaluating a loss function, denoted as $\epsilon_{\mathrm{FS}}$. Fig. 10 illustrates the behavior of the RAE (34), computed for three VQAs (cf. Section IV.2), including BENQO, as a function of the number of shots $n_{\mathrm{shots}}$ (at a fixed system size) and the system size $n$ (at a fixed number of shots). As expected from a multinomial probability distribution, the error decreases proportionally to $1/\sqrt{n_{\mathrm{shots}}}$ when $n$ is fixed. Conversely, when $n_{\mathrm{shots}}$ is kept constant, the error grows exponentially, reflecting the increasing range of loss values alongside a simultaneously shrinking variance as the system size increases. For an analytical derivation of this scaling behavior, see Appendix C (cf. [39]).
@@
The figure clearly illustrates that if the boundary line identified in Section V.2 was to follow an exponential decay, the number of required measurement shots would already grow faster than the number of brute-force trials needed to solve the problem classically. This strongly suggests that quantum advantage for optimization tasks under these conditions is fundamentally unattainable. In the alternative scenarios of power-law and logarithmic scaling, a small window of potential quantum advantage remains for problem sizes $n\geq 20$ (logarithmic boundary) and $n\geq 25$ (power-law boundary), where the shot requirement still falls below the critical threshold of quantum disadvantage. However, these potential advantages are only theoretical, as they are based purely on resource scaling and do not yet translate into practical runtime improvements.
@@
VII Validation and Discussion
A common concern in empirical studies like ours is to what degree the observed results are influenced by specific choices in the experimental setup (cf. Fig. 3). Factors such as problem formulation, quantum ansatz, noise model, optimizer selection, and performance metrics may, in principle, affect the conclusions drawn. One could, therefore, argue that the critical behavior identified in this work merely arises from unlucky experimental choices rather than inherent scalability challenges of hybrid variational quantum optimization. While each of our decisions was carefully justified to favor well-behaved scenarios and ensure a fair and optimistic evaluation, this section provides additional validation and discussion about the reliability and generality of our findings.
VII.1 Independent Confirmation of Scaling Trends
A study by Scriva et al. [18] examined shot noise effects using a VQE ansatz on an Ising chain problem and found that the success probability for a fixed system size follows the same sigmoidal pattern as observed in our experiments. Their results also demonstrated that the optimal number of function calls required to reach a fixed success probability scales exponentially with system size for certain parameter initializations, reinforcing our findings in Section VI.2.
@@ VII.1 Independent Confirmation of Scaling Trends
Despite the differences in quantum algorithms, noise models, classical optimizer selection, and problem types, both studies independently observed the same key trends as those presented in this work.
VII.2 Validation with a VQE Ansatz
We repeat our experiments to further validate our conclusions using another well-behaved quantum loss function: the two-local shallow VQE ansatz described in Section IV.2. The results for the NGD optimizer are summarized in Fig. 12, closely matching those obtained using BENQOs. This agreement is not only qualitative but also quantitative. Normalizing the noise resilience threshold $\sigma^{*}(n)$ by the variance of the respective loss functions (cf. Fig. 4) – thus mapping both results onto the RAE metric (34) – would likely reveal an even stronger correspondence between the two ansatzes.
Figure 12: Validation of NGD solvability scaling with VQE. (a) Contour plot of the empirical NGD solvability experiments (cf. Fig. 5), this time using VQE as the underlying quantum loss function. (b) Scaling of the success probability $\hat{p}_{\mathrm{opt}}$ across different noise levels for fixed system sizes, fitted with sigmoidal theory curves as in Fig. 6. Blue vertical lines indicate the points of steepest descent $\sigma^{*}$ (27), obtained from $tanh$-fits (cf. Eq. 25). (c) The main noise resilience parameter $\sigma^{*}$ is plotted against increasing system sizes and directly compared to the noise threshold obtained for BENQO (cf. Section V.2). The curves exhibit a close agreement.
To directly validate the corresponding resource scaling reported in Section VI.2, we conduct additional experiments with the VQE algorithm, this time introducing noise solely through finite sampling rather than the additive Gaussian noise model. The shallowness of the VQE ansatz enables simulations with larger system sizes than in previous experiments, extending the range to $n\in[3,13]$. The following results focus on the COBYLA solver, as it demonstrated the highest resource efficiency in terms of quantum function calls and mean runtime (cf. Appendix D).
@@ VII.2 Validation with a VQE Ansatz
VII.3 Reconsidering the Success Metric
Another point of discussion of our study is the choice of success metric. In this study, we measured the sample success probability of finding a solution within a specific relative threshold $t$ of the optimal energy (cf. Eq. 23), with thresholds set at $t=100\%$ (exact optimality), $99\%$, and $95\%$ (approximate solvability). However, most of the scalability issues discussed are derived from the probability of finding the optimal solution $\hat{p}_{\mathrm{opt}}$. One could argue that the primary goal of quantum optimization, particularly for NP-hard problems, is not necessarily to find the absolute global optimum but rather to generate high-quality approximations that outperform classical solutions.
Our experimental results indicate that the scaling behavior of $\hat{p}_{99\%}$ and $\hat{p}_{95\%}$ differs slightly from that of the exact solution probability (cf. Fig. 7). In particular, the further the relative energy threshold $t$ deviates from $100\%$, the more its scaling diverges from the solvability scaling discussed earlier. In some cases – especially in the absence of noise – the near-optimal success probability $\hat{p}_{95\%}$ even increases with system size (cf. Fig. 5), suggesting that as the system grows, optimizers tend to converge to local minima that lie within a close vicinity of the global optimum.
This trend can be explained by examining the distribution of loss values across all possible solution states (cf. Appendix E). In our QUBO formulation, feasible solutions are enforced via quadratic penalty terms (cf. Appendix B), which can – depending on the chosen penalty factor – significantly expand the loss value range in one direction, often leading to $|C_{\max}|\gg|C_{\min}|$. Due to the chosen normalization, the maximum loss value remains 1, while the majority of loss values concentrate (possibly exponentially) near zero (cf. Section IV.2), leading to an asymmetry in the loss value distribution. As the system size increases, this effect intensifies, effectively pushing most local minima closer to the global optimum. Consequently, the approximation ratio in Eq. 22 approaches 1, shifting the threshold for identifying feasible solutions to much higher energy levels – sometimes even beyond the $99\%$ threshold used in this study [104, 36]. This explains why approximate solutions remain relatively accessible, even when finding the exact optimum becomes increasingly difficult.
@@ VII.3 Reconsidering the Success Metric
VII.4 On Ansatz Choice and Smart Initializations
In this study, we deliberately focused on two variational ansatzes that avoid barren plateaus, maintaining non-vanishing gradient and loss values to ensure more favorable trainability properties on average (cf. Section IV.2). However, their absence of barren plateaus is ultimately linked to a non-exponentially scaling dimension of the underlying dynamical Lie algebra [26, 25], which in turn implies efficient classical simulability [11, 12]. As such, these ansatzes may not offer a viable path toward achieving quantum advantage within VQAs.
To address this, one might intentionally consider ansatz structures that do exhibit BPs, such as QAOA or generic hardware-efficient ansatzs (HEAs) [24], in order to retain the possibility of quantum advantage in the loss computation. In these cases, the key challenge becomes overcoming the exponential suppression of loss and gradient magnitudes. A leading approach to mitigating this issue is the use of smart parameter initialization strategies, which have shown promising performance in both QAOA [106, 107, 18, 108, 109, 110] and HEA circuits [111, 112, 86, 113, 114]. An important direction for future work is to investigate how the presence of Gaussian uncertainty interacts with such initialization strategies, and whether the observed impact of stochastic noise differs when applied to barren-plateau-prone ansatzes initialized with trainability-enhancing methods.
VIII Conclusions and Outlook
@@
The potential of variational quantum algorithms for optimization fundamentally depends on their ability to handle stochastic noise, which arises both from current hardware imperfections and finite sampling – the latter remaining even in fault-tolerant quantum computers. While numerous theoretical and empirical studies have highlighted how hardware noise can impair or even eliminate quantum advantage in VQAs (cf. Section II), there has been less focus on the impact of isolated stochastic uncertainty on system solvability. This work provides a rigorous numerical investigation into how classical optimizers – the main computational engine of VQAs – respond to additive Gaussian noise as the size of quantum loss functions scales up. Our findings, validated through extensive testing, empirically demonstrated that stochastic noise alone may impose fundamental limitations on the scalability of variational quantum optimization to practically relevant problem sizes, even when using fault-tolerant hardware.
@@
We find, that for fixed system sizes, the probability of finding high-quality solutions – including those within $1\%$ and $5\%$ of the optimal energy – decreases in a sigmoidal pattern with respect to the logarithmic noise level. This reveals a sharp performance drop at a critical noise threshold $\sigma^{*}$, which varies significantly across the different optimizers. Using $\tanh$-fits, we extracted these thresholds for system sizes $n\in[3,10]$, suggesting a common scaling trend across all optimizers: their noise resilience declines rapidly, potentially exponentially, as system size increases. Notably, this decline is steeper than what can be explained by a simultaneously shrinking loss variance, confirming that larger systems inherently present a more challenging loss landscape.
@@
While it is well known that current hardware error rates far exceed what is required for classically relevant problem sizes, our results highlight another possible limitation imposed by finite sampling errors. Extrapolating the observed decline in noise resilience showed that, to achieve the loss precision necessary for problem solvability, the required number of measurement shots scales prohibitively bad with system size. This leaves minimal room before exceeding the computational resources of classical brute-force sampling, which is highly impractical. These findings suggest that achieving meaningful quantum advantage with VQAs for large-scale optimization problems may face significant challenges.
@@
We therefore advocate for a shift in research focus – away from introducing yet new heuristic hybrid algorithms with idealized testing, and toward strategies that directly address the challenges demonstrated in this and prior work. If quantum utility [115] in optimization remains the goal, future research should prioritize alternative algorithmic frameworks (and possibly different problem domains) that can provably circumvent these barriers.
@@
Apart from the methodological suggestions made in Sections VII.3 and VII.4, one promising avenue is improved problem encoding, which aims to reduce the number of required qubits and thereby mitigate some of the scaling issues discussed here [116, 117, 118, 119, 120]. However, while these methods can offer potential efficiency gains, they often lack both theoretical performance guarantees and large-scale empirical validation in noise-exposed settings. Moreover, given that the scalability challenges demonstrated here may be inherent in the hybrid nature of VQAs, which rely on classical optimization, it seems to be more promising to explore fully quantum alternatives outside the variational paradigm. Recent works [121, 122, 123, 120] have begun such investigations, and a similar shift away from hybrid frameworks can be seen in the quantum machine learning community [124]. Additionally, this study has not yet considered alternative computational models, such as Quantum Boson Sampling [125, 126, 127] or Quantum Annealing [128, 129], which may warrant further investigation.
@@
Ultimately, our work provides a general and robust framework for rigorously investigating the scalability and limitations of not only VQAs but also other quantum algorithms in the presence of stochastic errors. We encourage researchers to apply and extend our methodology to more promising alternative setups – whether by testing different quantum ansatzes, optimizer setups, problem instances, or hardware platforms. If future studies uncover scenarios where problem solvability scales more favorably with system size and noise levels, this would be an exciting breakthrough. However, if the scaling behavior observed here proves to be universal and inherent to the classical optimization problem inherent in hybrid architectures like VQAs, it would strongly suggest that quantum advantage in hybrid variational quantum optimization is not realistic. In fact, our results indicate that the challenges facing VQAs – both in the NISQ and fault-tolerant era – may be more fundamental than what can be resolved through hardware improvements, better classical optimizers, or ansatz tuning alone.